a clothing store offers a 20% discount on all items of the store. If the original price of a pair of jeans is 84.00 how much will it be after the discount





Answer by stanbon(75887)

a clothing store offers a 20% discount on all items of the store. If the original price of a pair of jeans is 84.00 how much will it be after the discount
----------------
Equation:
price after discount = 84-0.20*84 = 0.80*84 = $67.20
----------------------
Cheers,
Stan H.
--------------